OpenStudy (anonymous):

lim X^4 -16X / X^2 + 2X x -> 0 Help please!

OpenStudy (anonymous):

lim x -> 0 \[\frac{ X^4 - 16X }{ X^2 +2X }\]

jimthompson5910 (jim_thompson5910):

try to factor each piece as much as possible

jimthompson5910 (jim_thompson5910):

you'll find that something cancels

OpenStudy (anonymous):

\[X(X^3 - 16)\] Is this correct?

OpenStudy (anonymous):

Like for the top?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

\[X(X+2)\] = \[\frac{ X(X^3 - 16 )}{ X(X-2) }\]

jimthompson5910 (jim_thompson5910):

x+2 not x-2

jimthompson5910 (jim_thompson5910):

but you have the right idea

OpenStudy (anonymous):

yeah sorry wrong button.

jimthompson5910 (jim_thompson5910):

what terms cancel?

OpenStudy (anonymous):

oh god. I think I figured it out

OpenStudy (anonymous):

X's cancel

jimthompson5910 (jim_thompson5910):

yes, the ones to the left

OpenStudy (anonymous):

Then change the other X's to 0 and solve

jimthompson5910 (jim_thompson5910):

yes, then plug in x = 0 and you no longer have to deal with a division by zero error

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome